NSF Lube is a fast-growing chain of oil-change stores. The following data are available for last year’s services.

  • NSF Lube performed 468,200 oil changes last year. It had budgeted 434,800 oil changes, averaging 12 minutes each.
  • Standard variable labor and support costs per oil change were as follows.
Direct oil specialist services: 12 minutes at $23 per hour $ 4.60
Variable support staff and overhead: 12.0 minutes at $17 per hour 3.40
  • Fixed overhead costs:

Annual budget $1,038,800

  • Fixed overhead is applied at the rate of $2.40 per oil change.
  • Actual oil change costs:
Direct oil specialist services: 468,200 changes averaging 21 minutes at $26 per hour $ 4,260,620
Variable support staff and overhead: 0.21 labor-hours at $16 per hour × 468,200 changes 1,573,152
Fixed overhead 1,410,000

Required:

a. Prepare a cost variance analysis for each variable cost for last year.
b. Prepare a fixed overhead cost variance analysis.

omplete this question by entering your answers in the tabs below.

  • Required A
  • Required B

Prepare a cost variance analysis for each variable cost for last year. (Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)

Price Variance Efficiency Variance Total Variance
Oil specialist    U    U U
Variable overhead F U F

Complete this question by entering your answers in the tabs below.

  • Required A
  • Required B

Prepare a fixed overhead cost variance analysis. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect.)

Price variance U
Production volume variance F
Fixed overhead cost variance U
Answer #1
Standard ( 1 oil change) Standard ( 468200 oil changes) Actual (468200 oil changes)
Hours Rate Amount Hours Rate Amount Hours Rate Amount
Oil specialists 0.2 $           23.00 $      4.60 93640 $    23.00 $ 2,153,720.00 163870 $    26.00 $ 4,260,620.00
Variable support staff 0.2 $           17.00 $      3.40 93640 $    17.00 $ 1,591,880.00 98322 $    16.00 $ 1,573,152.00
Price Variance = (Standard Price - Actual Price) x Actual Hours Efficiency Variance = (Standard Hours - Actual Hours) x Standard Price Total Variance = Standard Cost - Actual Cost
Oil Specialists = ($ 23 - $ 26) x 163870 Oil Specialists = (93640 - 163870) x $ 23 Oil Specialists = $ 2153720 - $ 4260620
= $ 491,610.00 (Unfavorable) = $1,615,290.00 (Unfavorable) = $   2,106,900.00 (Unfavorable)
V.Support staff = ($ 17 - $ 16) x 98322 V.Support staff = (93640 - 98322) x $ 17 V.Support staff = $ 1591880 - $ 1573152
= $   98,322.00 (Favorable) = $      79,594.00 (Unfavorable) = $         18,728.00 (Favorable)
Fixed OH
Budgeted OH $ 1,038,800.00
Recovery Rate $                2.40 per oil change
Actual no. of oil changes 468200
Recovered OH $ 1,123,680.00
Actual OH $ 1,410,000.00
Fixed OH price variance = Budgeted OH - Actual OH
= $ 1038800 - $ 1410000
= $    371,200.00 (Unfavorable)
Fixed OH Volume Variance = Recovered OH - Budgeted OH
= $ 1123680 - $ 1038800
= $      84,880.00 (Favorable)
Fixed OH cost variance = Recovered OH - Actual OH
= $ 1123680 - $ 1410000
= $    286,320.00 (Unfavorable)